### v2

# Descriptive Caption Enhancement with Visual Specialists for Multimodal Perception

Abstract
Training Large Multimodality Models (LMMs) relies on descriptive image caption that connects image and language. Existing methods either distill the caption from the LMM models or construct the captions from the internet images or by human. We propose to leverage off-the-shelf visual specialists, which were trained from annotated images initially not for image captioning, for enhancing the image caption.
Our approach, named DCE, explores object low-level and fine-grained attributes (e.g., depth, emotion and fine-grained categories) and object relations (e.g., relative location and human-object-interaction (HOI)), and combine the attributes into the descriptive caption. Experiments demonstrate that such visual specialists are able to improve the performance for visual understanding tasks as well as reasoning that benefits from more accurate visual understanding. We will release the source code and the pipeline so that other visual specialists are easily combined into the pipeline. The complete source code of DCE pipeline and datasets will be available at https://github.com/syp2ysy/DCE.
1 Introduction
Recent advancements in Large multimodal models (LMMs) [53, 55, 60] have significantly enhanced the understanding and reasoning abilities for multimodal tasks. Vision-language connection is crucial for high abilities, and descriptive image captions serve as one of key components for image perception. The image caption is expected to describe the image as detailed and complete as possible. There are two main categories for image captions. One is to generate image captions from human annotation, such as COCO [30]and LAION [43]. However, the high cost of human annotation limits scalability. The other one is LMMs [11, 1] annotation, such as ShareGPT4V [8] and Densefusion [26]. While captions generated by LMMs offer better scalability, their comprehensiveness and accuracy often fall short.
There is still much improvement space for the captions from SoTA LMMs, such as InternVL2 [11] and from current human annotation datasets. An example is given in Figure 1(a) for showing the improvement space. It can be seen that the caption from COCO [29] a widely-used human-annotated dataset, is usually incomplete, only describes a small portion of the image content, and a lot of information is missing. The captions from LMMs are much better, but still can be improved in some aspects. As shown in Figure 1(b), we selected 8 objects and 5 key attributes from the image to analyze and compare the captions generated by different methods. The key attributes include fine-grained attributes, spatial relation, and HOI. It is clear that captions generated by LMMs are more detailed than those annotated by humans. They not only describe more objects but also more attributes. However, captions from LMMs still much improvement space, as they tend to overlook important objects and attributes. For example, in Figure 1(a), Object 6 is completely ignored by all LMMs. Furthermore, crucial 3D spatial relationships between objects are also missing, which is particularly problematic for tasks that require a comprehensive understanding of the scene’s structure [21, 34].
Figure 1: (a) We present a comparison of captions from DCE, human, and generalist LMM models annotations, including InternVL2-26B, LLaVA-NeXT, and GPT-4V. (b) visualizes the extent to which the captions in (a) describe multiple objects and various attributes, including Objects 1-8, Object Attributes, OCR, HOI, 2D spatial relations and 3D spatial relations.
Toward this end, we propose a Descriptive Caption Enhancement Engine (DCE), designed to enable efficient and low cost image captioning. We leverage visual specialists [59, 9, 48, 54] to replicate various human visual capabilities, and subsequently employ large language models (LLMs) [51, 4] to simulate the human cognitive process. This combined approach enables us to generate high-quality image captions by closely mimicking the way humans perceive and interpret visual information. Notably, DCE relies solely on open-source visual expert models and LLMs, significantly reducing annotation costs.
Specifically, we leverage existing visual specialists to obtain instance-level and relational attributes within images. Instance-level attributes focus on object low-level and fine-grained attributes (e.g., depth, emotion and fine-grained categories). Relational-level attributes capture interactions and relationships between objects (e.g., relative location and HoI). Next, we use prompts to guide LLMs in combining object attributes into region captions. Finally, prompts are used again to integrate these region captions with relational-level attributes, producing a comprehensive and detailed image caption. Since DCE utilizes multiple off-the-shelf visual specialists, the resulting captions capture a wide array of detailed attributes and nuanced relationships, leading to richer and more precise image descriptions. As shown in Figure 1(b), captions annotated by DCE contain the most comprehensive information among all methods. Figure 2 quantitatively demonstrates that the captions generated by DCE provide greatest benefit to LMMs.
We applied our DCE to annotate a large-scale dataset of 1.1 million images, consisting of 1 million diverse images (DCE-1M) and 118K real-world scene images (DCE-118K). Experiments were conducted using both LLaVA-v1.5 and LLaVA-NeXT models. The results show that the highly detailed captions generated by DCE significantly enhance the perceptual capabilities of large multimodal models (LMMs), improving visual-language alignment. The outstanding performance of both LLaVA-v1.5 and LLaVA-NeXT across 14 benchmarks further underscores the effectiveness of our approach, confirming that the generated image captions are of exceptional detail and quality.
Figure 2: Comparisons of caption quality. (a) and (b) show the downstream task performance of LLaVA-v1.5 and LLaVA-NeXT after pretraining with different image captions.
2 Related Work
Figure 3: The DCE pipeline first utilizes various visual specialists to extract both Object and Relation attributes. Then, it uses an LLM to integrate the object attributes into detailed region captions, followed by combining the region captions with relational attributes to generate a comprehensive image caption.
Table 1: Summary of attributes our approach extracts through visual specialists. It includes the specific attribute names, the models used, and the extraction process for each.
Attributes Visual Specialists Detailed Process
Object
Size Detection model Using the area of the bounding box to measure the size of the instance.
Depth Depth & Detection model Average the depth map values within the bounding box region to obtain the depth information.
Emotion Emotion model If the detected region is labeled as ”person”, an emotion model is used to extract an emotion label.
OCR OCR Model Using an OCR model to extract the text content and bounding box from the region.
Animal Fine Grained model A fine-grained recognition model to identify specific species of the animal.
Plants A fine-grained recognition model to identify specific species of the plants.
Aircrafts A fine-grained recognition model to identify specific model of the aircraft.
Logo A fine-grained recognition model to identify logos in the region.
Landmark A fine-grained recognition model to identify landmarks within the region.
Food A fine-grained recognition model to identify specific species of the food.
Celebrity Using a fine-grained recognition model to identify celebrity within the region.
Relation
P2O relation HOI Model
Using an HOI model to determine the relationship between the person and the object, while
the bounding boxes of both the person and the object define their respective regions.
Count Detection model Counting the number of all objects in the image based on the detection results.
2D Absolute Location Detection model
Using the bounding box to determine the instance’s position within the image, including regions
such as left, right, top, bottom, center, top-left, bottom-left, top-right, and bottom-right.
2D Relative Location Detection model
Using the bounding box to determine the relative position among multiple objects within
the image, including regions such as left, right, near, next to, close by, and so on.
3D Relative Location Detection & Depth model
Using the depth attributes of different instances to capture the 3D spatial relationships of objects
relative to the camera, such as ”Instance_A is in front of Instance_B” or ”Instance_A is behind of
Instance_B” relative to the camera.
Large multi-modality models. There are recently a lot of developments for large multi-modality models. One major effort lies in aligning/connecting the pretrained vision encoder and the the pretrained language model [53, 55, 50]. Flamingo [2] inserts new gated cross-attention layers between existing pretrained and frozen LM layers, that bridges powerful pretrained vision-only and language-only models. BLIP-2 [25] bridges the two modalities with a lightweight querying transformer. Qwen-VL [4] adopt a way that is similar to query transformer. LLaVA [33] adopt a simple projection model to connect the visual encoder and the language model. The Llama3 Herd of Models [14] uses a adapter similar to LLaVA and BLIP-2 to connect the vision encoder and the language model. Gemini [16] simply concatenates multi-modality tokens, such as image and text tokens, and feed them into a transformer as the input. Emu [49] uses a causal transformer to train the image encodings to tokens. There is also some effort lying in the encoder architecture for efficient image encoding. One approach involves dividing the image into large blocks, which helps capture finer-grained features while maintaining computational efficiency, such as Monkey [28], Qwen2-VL [52], LLaVA-NeXT [24]. Additionally, many works have further enhanced LMMs by focusing on the quality and diversity of pretraining and finetuning data [4, 40, 22, 7, 6, 53]. Despite the success of these LMMs, few studies have focused on obtaining large-scale high-quality image-text data, which is a crucial factor in driving the capabilities of LMMs.
Figure 4: The prompt for using LLM to generate an region caption by considering object attributes and reference captions.
Descriptive captions. CC3M [45] (Conceptual Captions: A Cleaned, Hypernymed, Image Alt-text Dataset For Automatic Image Captioning) is harvested from the Alt-text HTML attribute associated with web images followed by an automatic pipeline that extracts, filters, and transforms candidate image/caption pairs. CC12M [5] extends CC3M [45] with an emphasis on long-tail visual recognition. It builds on the same foundational methodology as CC3M but incorporates a more diverse range of objects, scenes, and rare visual concepts to better represent the breadth of real-world imagery. SBU Captions [39] forms the data with images and descriptions sourced from Flickr. The captions are manually written by users and describe a wide variety of visual content, from everyday scenes to more specialized imagery. COCO-captions [10] is built by collecting captions on Flickr images that are generated by human subjects on Amazon’s Mechanical Turk (AMT). While manually annotated captions effectively capture the main information in images, they often overlook important details and contextual richness. To address this, methods like ShareGPT4V [8] and Densefusion [26] leverage LLMs [8, 52] to generate more detailed captions. However, these LMM-based approaches still face challenges, particularly with over-simplified or inaccurate descriptions, which can compromise caption quality and reliability [41, 21]. Balancing detail richness and accuracy remains a key challenge in current research. DenseFusion and our DCE share a similar goal of generating more accurate image captions by integrating diverse visual information. However, DenseFusion relies on GPT-4V, a highly expensive model, whereas our approach significantly reduces costs and improves efficiency by utilizing open-source visual expert models and LLMs.
3 Approach
As shown in Figure 3, DCE leverage existing visual specialists to extract visual properties for improving descriptive captions. We explore two kinds of properties: object-level attributes and object-relations. Our approach consists of: object localization using a SoTA open-world object detection specialist, object property extraction using various specialists, and object relation extraction between objects.
3.1 Object Attributes
Object localization. We combine in-domain detection models [9, 38] and open-world detection models [18] for robust detection, merging bounding boxes from both models with confidence scores above 0.5. The object detection model outputs the location and semantic information of objects that existed in the image. After detecting these regions, we apply Non-Maximum Suppression (NMS) to eliminate redundant or overlapping boxes. The IoU threshold for NMS is $0.75$.
Figure 5: The prompt for LLM to generate an image caption by considering relation attributes, region location information and captions.
Attribute extraction by visual specialists. Table 1 presents all the instance-level attributes involved in DCE, along with their extraction processes. The attributes currently include three main parts. (1) Fine-grained object category attributes. This is rarely explored in current multimodal models [8]. In DCE, we incorporate fine-grained details by introducing various specialized models, covering categories such as animals, plants, food, logo, aircraft, landmarks, and celebrities. The animal and plant attributes contain the fine-grained category of $891k$ and $427k$ species of animals and plants, respectively. Each species within this category possesses unique characteristics and behaviors, making it a rich and varied classification. The food attributes include a variety of food types commonly found in daily life, while logos are graphic symbols that serve as visual identifiers for conveying messages. Landmarks are the famous tourist attractions that hold cultural, historical, or geographical significance. Aircraft refers to the machines designed for flight, including airplanes, helicopters, drones, and other aerial vehicles. Celebrities are individuals widely recognized in public life, entertainment, sports, or other fields. All this specific and fine-grained information could be regarded as the external world knowledge, aligning the textual content in the image with basic human cognition. (2) Low-level and emotion attributes. It includes emotion, depth, and size. (3) OCR. It is one of the important attributes for multimodal models. The specific visual expert models we use will be presented in the Appendix.
3.2 Object Relation
DCE can extract relationships between multiple objects. We consider three categories: the interactions between humans and objects, the 2D as well as 3D relative positional relationships among different objects, and the object counting information. Table 1 presents the relation attributes and their extraction process.
The human-object interaction provides essential information about the actions and activities performed by humans with the objects. We utilize the human-object interaction (HOI) model to detect interactive activities between humans and objects in the image. The interactions detected by the HOI model can be used to supplement events not mentioned in the caption.
Table 2: Human evaluation of attribute richness, conducted on 100 validation samples with 10 volunteers.
Attributes InternVL2 LLaVA-NeXT DCE
Spatial Relation 0.57 0.62 0.75
HOI 0.92 0.86 0.92
Fine-Grained 0.16 0.08 0.24
OCR 0.26 0.33 0.48
Emotion 0.23 0.14 0.47
Location 0.36 0.59 0.81
The 2D positional information captures the spatial relationships of objects, comprising both 2D Absolute Location and 2D Relative Location. The 2D Absolute Location describes an object’s position relative to the image (e.g., Object A is on the left side of the image), while the 2D Relative Location describes positional relationships between objects (e.g., Object A is next to Object B). We use the bounding boxes of objects to determine their positional relationships.
The 3D relative positional information captures the spatial relationships of objects in 3D space, defining both their absolute positions in the scene and their relative positioning (e.g., Object A is in front of Object B or at a specific angle). This information enhances the understanding of a scene’s 3D structure and provides richer spatial awareness. We leverage the depth differences between objects to determine their 3D positional relationships.
Table 3: Performance on seven General Visual Question Answering benchmarks. The red and blue colors respectively represent the optimal and suboptimal results on each benchmark. * indicates the use of LLaVA-NeXT’s open-source SFT data, with certain private data excluded.
Model LLM Visual Question Answering Benchmarks
VQAv2 DocVQA OKVQA GQA TextVQA ScienceQA Ai2d
Low Resolution Models
BLIP2 [25] Flan-T5 41.0 - 45.9 41.0 42.5 61.0 -
InstructBLIP [13] Vicuna-7B - - - 49.2 50.1 60.5 40.6
InstructBLIP [13] Vicuna-13B - - - 49.5 50.7 63.1 -
IDEFICS-Instruct [20] LLaMA-65B 60.0 - 36.9 - 32.9 61.8 54.8
OpenFlamingo [3] MPT-7B 53.0 - 38.3 - 28.3 44.8 -
InternVL-Chat [11] Vicuna-7B 79.3 - 51.8 62.9 57.0 - -
Qwen-VL-Chat [4] Qwen-7B 78.2 62.6 56.6 57.5 61.5 68.2 -
mPLUG-Owl2 [55] LLaMA-7B 79.4 - 57.7 56.1 58.2 68.7 55.7
LLaVA-v1.5 [31] Vicuna-7B 78.5 28.1 - 62.0 58.2 66.8 55.5
ShareGPT4V [8] Vicuna-7B 80.6 - - 63.3 60.4 68.4 -
LLaVA-v1.5(Ours) Vicuna-7B 80.9 39.1 57.2 64.2 61.4 71.0 59.4
High Resolution Models
Monkey [28] Qwen-7B 80.3 66.5 61.3 60.7 67.6 69.4 62.6
LLaVA-NeXT [24] Vicuna-7B 81.8 74.4 44.3 64.2 64.9 70.1 66.6
LLaVA-S2 [46] Vicuna-7B 79.7 - - 63.3 60.8 68.2 -
LLaVA-HR [36] Vicuna-7B 81.9 - 58.9 64.2 67.1 65.1 -
LLaVA-NeXT(Ours) Vicuna-7B 82.4 78.8 57.2 65.2 64.8 71.2 71.2
3.3 Captioning with Attributes
Region captioning. We use a large language model (Qwen-72B) to integrate the object attributes with the caption from a large multimodal model (InternVL2-26B). For example, in cases where the fine-grained model for “animals” identifies a specific class label like “$\{animal\_name\}$, we employ the prompt $\{cat\_name\}$ exists in the region and $\{animal\_name\}$ is the $\{cat\_name\}$’s subclass; use $\{animal\_name\}$ in the caption; otherwise, do not mention $\{animal\_name\}$ to guide the LLM in integrating this specific label into the caption. In this process, $\{cat\_name\}$ represents the coarse-grained label provided by the detection model. The detailed prompt engineering process is illustrated in Figure 4.
Image captioning. We combine object relation attributes, and region location information (for object grounding) and captions to get improved image captions. We use LLM for the combination. The prompt is given in Figure 5. For example, to describe the 3D relative positional relationship between $\{bbox\_0\}$ and $\{bbox\_1\}$, we use the prompt: ”Relative to the camera, the $\{category\_0\}$ in $\{bbox\_0\}$ in the image is $\{3d\_relation\}$$\{category\_1\}$ in $\{bbox\_1\}$, where $\{3d\_relation\}$ can be ”in front of” or ”behind of”. We then use a large language model (Qwen2-72B-AWQ) to integrate this information with the detailed region caption and the image reference caption.
3.4 Analysis
Dataset Description. We leverages DCE to generate more detailed annotations for publicly available image datasets, resulting in two enhanced datasets: (1) DCE-1M, which provides dense annotations for 1 million diverse images from Densefusion [26], covering a wide range of objects and scenes, and (2) DCE-118K, which includes refined annotations for 118,000 complex scene images from the COCO dataset [30]. We provide a detailed analysis of the DCE-annotated image captions in the Appendix.
Attribute richness. We randomly select 100 images from the DCE-118K and invite five independent evaluators to assess the captions for each image. The evaluators analyzed and recorded the occurrence of the attributes within the captions, such as spatial relations, HOI, and OCR (optical character recognition) details.
The results as show in Table 2. Compared to other models and human annotations, our approach demonstrated a greater ability to capture and express a wider diversity of visual attributes present in the images. This suggests that our approach provided richer and more detailed descriptions of the visual content.
Comparison. We conducted a comparison of human annotations, internVL2-26B, LLaVA-NeXT-34B, and DCE-annotated 118K datasets on both LLaVA-v1.5 and LLaVA-NeXT. The results as shown in Figure 2, the captions annotated by DCE significantly improve the performance of LLaVA-v1.5 and LLaVA-NeXT. Compared to human and generalist multimodal model annotations, DCE-annotated captions are richer and more detailed, offering deeper context and capturing finer nuances. This enhanced descriptive quality leads to better performance in downstream tasks.
4 Experiments
In this section, we first present the implementation details, then compare the DCE-1M-trained model with state-of-the-art LMMs across multiple benchmarks. Finally, we perform ablation studies on the DCE-118K-trained model to validate the effectiveness of DCE.
Table 4: Performance on seven Large Multi-Modal benchmarks. The red and blue colors respectively represent the optimal and suboptimal results on each benchmark. * indicates the use of LLaVA-NeXT’s open-source SFT data, with certain private data excluded.
Method Vision Encoder Language Model MMBench-CN MMBench MM-Vet SEEDI SEED-Bench MMMU POPE
Low Resolution Models
BLIP-2 [25] ViT-g (1.3B) Vicuna-7B - - 22.4 46.4 - - 85.3
MiniGPT-4 [58] ViT-g (1.3B) Vicuna-7B 11.9 23.0 22.1 - 47.4 23.6 -
InstructBLIP ViT-g (1.3B) Vicuna-7B 23.7 36.0 26.2 53.4 - 30.6 78.9
LLaMA-Adapter-v2 [14] ViT-L (0.3B) LLaMA-7B - 39.5 31.4 - 32.7 - -
OpenFlamingo [3] ViT-L (0.3B) MPT-7B - 5.7 24.8 - 42.7 26.3 -
Otter [22] ViT-L (0.3B) LLaMA-7B - 48.3 24.6 - 32.9 - -
Qwen-VL-Chat [4] ViT-G (1.9B) Qwen-7B 56.7 60.6 - 58.2 - 29.6 -
mPLUG-Owl2 [55] ViT-L (0.3B) Vicuna-7B - 64.5 36.2 - 57.8 34.7 86.2
LLaVA-v1.5 [31] ViT-L (0.3B) Vicuna-7B 57.6 64.3 30.5 66.2 58.6 35.3 85.9
ShareGPT4V [8] ViT-L (0.3B) Vicuna-7B 62.2 68.8 37.6 69.7 61.9 - 85.7
LLaVA-v1.5(Ours) ViT-L (0.3B) Vicuna-7B 60.0 69.2 38.2 70.3 64.3 36.3 86.4
High Resolution Models
LLaVA-NeXT [24] ViT-L (0.3B) Vicuna-7B 60.6 67.4 43.9 70.2 64.7 35.1 86.5
ShareGPT4V-S2 [26] ViT-L (0.3B) Vicuna-7B - 68.0 35.0 70.1 62.4 - 86.7
LLaVA-S2 [46] ViT-L (0.3B) Vicuna-7B - 66.4 34.6 67.2 59.9 - 86.7
LLaVA-HR [36] ViT-L (0.3B) Vicuna-7B - - 31.2 - 64.2 - 87.6
LLaVA-NeXT(Ours) ViT-L (0.3B) Vicuna-7B 61.7 69.3 40.1 72.2 65.7 36.0 87.0
4.1 implementation details
Model and Training Set. We conduct experiments on LLaVA-v1.5 [31] and LLaVA-NeXT [32] to demonstrate the effectiveness of DCE. Specifically, we using CLIP-L [42] as the visual encoder and Vicuna-v1.5 [12] as the large language model. We adopt a two-stage training strategy: (1) Pre-Training Stage. We train only the projector for initial alignment. Then, following SharGPT4V [8], we set the last 12 layers of the visual encoder in LLaVA-v1.5 as trainable and make the entire LLaVA-NeXT model trainable, following [23], to further enhance perceptual capabilities. (2) Instruction Tuning Stage. We use the open-source LLaVA-mix-665K and LLaVA-NeXT-data to respectively train the LLaVA-v1.5 and LLaVA-NeXT models. The detailed training procedure is provided in the supplementary material.
Evaluation Benchmarks. We evaluate on seven visual question answering (VQA) tasks across domains such as document understanding, general knowledge, and scientific reasoning, including VQAv2 [15], DocVQA [37], OKVQA [44], GQA [17], TextVQA [47], ScienceQA [35], and Ai2d [19]. Additionally, we evaluate performance on five widely used LMM benchmarks designed to test multimodal models on visual grounding, scene understanding, and generalization: MMBench [34], MM-Vet [56], SEED [21], MMMU [57], and POPE [27].
Comparison Method. We compare LLaVA-v1.5 and LLaVA-NeXT models trained on DCE-1M against current SOTA MLLMs. Additionally, we evaluate the performance of various MLLMs in generating image captions, using the advanced models InternVL2-26B [11] and LLaVA-NeXT-34B [24] to generate competitive captions on the DCE-118K images.
4.2 Main Results
VQA Benchmarks. The results on six common visual question answering (VQA) datasets are presented in Table 3. It is clear that LLaVA-v1.5 and LLaVA-NeXT, trained with DCE-1M, achieve state-of-the-art performance in both low-resolution and high-resolution settings. Compared to the baseline LLaVA-v1.5 [31], our model excels across all VQA benchmarks, demonstrating that high-quality image captions significantly enhance model performance. This highlights the crucial role of detailed and accurate captions in improving visual understanding for VQA tasks. Furthermore, when compared to the baseline LLaVA-NeXT [24], this improvement remains consistent, suggesting that the impact of high-quality captions is not dependent on the model variation. Additionally, compared to models like ShareGPT-4V, our model demonstrates superior performance across most VQA benchmarks. This improvement indicates that the captions generated by our DCE method provide richer and more comprehensive information.
Table 5: Comparison of Different Image Captioning Annotation Methods.
      Annotation       OKVQA       GQA       ScienceQA       TextVQA       MMBench       MM-Vet       SEED-Bench
      LLaVA-v1.5
      + human [29]       54.9       62.4       68.6       58.1       65.0       31.6       61.1
      + InternVL2-26B [11]       54.7       63.0       69.1       58.4       64.8       32.7       61.8
      + LLaVA-NeXT-34B [24]       55.7       62.9       68.8       58.7       65.3       33.0       61.7
      + DCE       56.9       63.2       69.8       58.9       66.6       33.9       62.0
      LLaVA-NeXT
      + InternVL2-26B [11]       54.3       65.1       70.1       61.2       66.7       37.3       64.7
      + LLaVA-NeXT-34B [24]       54.3       65.0       70.5       61.0       67.2       37.2       64.5
      + DCE       56.7       65.2       72.0       62.0       68.5       37.8       65.0
Figure 6: Visualization of DCE’s Attribute Fusion: DCE combines object and relational attributes to generate detailed and comprehensive captions.
The model trained on DCE-1M demonstrates exceptional performance on datasets such as VQAv2 [15] and Ai2D [19], highlighting that integrating detection models into DCE significantly enriches the diversity of objects in the generated captions, thereby boosting the model’s performance on object recognition benchmarks. Furthermore, incorporating relational attributes into DCE enriches the captions with detailed inter-object relationships, enhancing the model’s ability to capture and understand complex relationships, which further improves its performance on visual reasoning benchmarks such as GQA [17]. However, performance on tasks like TextVQA [47] is hindered by limitations in the open-source OCR model and the threshold settings; a high threshold restricts the model’s ability to capture finer textual details.
Large Multi-Modal Benchmarks. We further conduct the evaluation on five challenging large multi-modal benchmarks. The experimental results are shown in Table 4. It can be seen that both LLaVA-v1.5 and LLaVA-NeXT trained with DCE-1M achieve competitive performance on more complex LMM benchmarks, demonstrating that the improvements brought by DCE-1M are comprehensive. Our model outperforms both LLaVA-v1.5 [31] and LLaVA-NeXT [32] across all LMM benchmarks, demonstrating that high-quality image captions during pretraining significantly enhance model performance, even without altering the supervised fine-tuning (SFT) data. Compared to other image captioning methods, such as ShareGPT-4V [8], DCE-generated captions provide richer and more comprehensive scene information, significantly boosting model performance across most LMM benchmarks. However, due to the lack of Chinese data in DCE-1M, the model performs poorly on MMBench-CN [34]. This highlights the need for multilingual image captioning, which will be an area for future improvement in DCE. Additionally, the detection model in DCE introduces some noise, which may interfere with the model’s ability to accurately capture objects, leading to decreased performance on tasks like POPE [27]. Therefore, reducing this noise will be a key focus for future improvements in DCE.
4.3 Ablation Study
Comparing different annotation methods. We compared different image annotation methods, including human annotations, GenerateList LMM annotations, and our DCE. Specifically, we annotated 118K COCO images and conducted comparisons on LLaVA-v1.5 and LLaVA-NeXT. The experimental results are shown in Table 5. We found that the image captions generated by DCE improve LMM performance on downstream tasks more effectively than other annotation methods. Notably, compared to captions annotated by internVL2, DCE’s inclusion of object attributes significantly improves model performance on OKVQA and TextVQA tasks. Relational attributes enhance the model’s understanding of multi-object relationships, leading to a notable increase in GQA performance. Furthermore, superior results on complex LMM benchmarks like MM-Vet, MMBench, and SEED highlight that DCE-generated captions provide rich and comprehensive scene information.
Case Study. Figure 6 presents example captions generated by the DCE engine and the general MLLM LLaVA-NeXT 34B. It is evident that visual specialists within DCE capture detailed object and relational attributes, resulting in richer and more descriptive captions. For instance, in Figure 6(a), the OCR information highlights the precision of OCR specialist model, while the fine-grained model’s identification of specific aircraft types further enhances caption informativeness. Additionally, the relational attributes significantly enrich description by providing detailed spatial relationships between objects, underscoring the advantages of DCE in capturing comprehensive scene details. More visualization results are provided in the Appendix.
5 Conclusion
In this paper, we propose a new image captioning engine, DCE , which utilizes off-the-shelf visual specialists to enhance the quality and detail of image captions. Unlike existing methods that primarily rely on LMM models or human annotation, DCE leverages visual specialists to simulate human perceptual abilities, while using LLMs to emulate cognitive processes. This dual approach enables DCE to generate captions that are both visually detailed and contextually aware. Through extensive experiments, we demonstrate that incorporating these visual specialists leads to improved model performance across various visual understanding and reasoning tasks, particularly those that depend on accurate attribute and relationship recognition. Our approach highlights the potential of leveraging specialized visual features for enhancing multimodal representations and provides a flexible framework for future integration of additional visual expertise. We will release the DCE source code and pipeline to facilitate further research, enabling the community to easily integrate other visual specialists and extend the capabilities of multimodal models.